OpenStudy (anonymous):

Which equation would best help solve this problem? Five added to 2 times a number is equal to 7 less than 4 times the number. A. 2x + 5 = 4x – 7 B. 4x +7 = 2x – 5 C. 5x + 2 = 7x – 4 D. 2x – 5 = 4x + 7

OpenStudy (anonymous):

I think its A

OpenStudy (anonymous):

@Vocaloid can you tell me if I'm right?

Vocaloid (vocaloid):

yes, good!

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

it c

OpenStudy (anonymous):

No its A

OpenStudy (anonymous):

Its A

OpenStudy (anonymous):

Because x=6

OpenStudy (anonymous):

2*6=12 12+5 =17 4*6=24-7=17

OpenStudy (anonymous):

srry I just had to re -read it its A

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I tend to accidentally read to fast